# CReFT-CAD: Boosting Orthographic Projection Reasoning for CAD via Reinforcement Fine-Tuning

### v1

@@ Abstract
Computer-Aided Design (CAD) plays a pivotal role in industrial manufacturing. Orthographic projection reasoning underpins the entire CAD workflow, encompassing design, manufacturing, and simulation. However, prevailing deep-learning approaches employ standard 3D reconstruction pipelines as an alternative, which often introduce imprecise dimensions and limit the parametric editability required for CAD workflows. Recently, some researchers adopt vision–language models (VLMs), particularly supervised fine-tuning (SFT), to tackle CAD-related challenges. SFT shows promise but often devolves into pattern memorization, yielding poor out-of-distribution performance on complex reasoning tasks. To address these gaps, we introduce CReFT-CAD, a two-stage fine-tuning paradigm that first employs a curriculum-driven reinforcement learning stage with difficulty-aware rewards to build reasoning ability steadily, and then applies supervised post-tuning to hone instruction following and semantic extraction. Complementing this, we release TriView2CAD, the first large-scale, open-source benchmark for orthographic projection reasoning, comprising 200,000 synthetic and 3,000 real-world orthographic projections with precise dimension annotations and six interoperable data modalities. We benchmark leading VLMs on orthographic projection reasoning and demonstrate that CReFT-CAD substantially improves reasoning accuracy and out-of-distribution generalizability in real-world scenarios, offering valuable insights for advancing CAD reasoning research. The code and adopted datasets are available at https://github.com/KeNiu042/CReFT-CAD.
@@ 1 Introduction
Computer-Aided Design (CAD) is now integral to industrial product development, driving design, manufacturing, and simulation workflows. In the design phase, engineers use CAD drawings or rasterized orthographic projections for their precision and easy editability. During manufacturing, these drawings are converted into constraint-based parameter tables, and for simulation, they yield boundary-representation (B-Rep) data or textual geometry descriptions. A truly user-centric pipeline thus requires accurate semantic parsing of orthographic projection: automated extraction of parameter tables directly produces 3D models that meet stringent manufacturing and simulation standards. By contrast, existing reverse-engineering methods rely on expensive 2D/3D scanning hardware and labor-intensive post-processing, severely limiting scalability and broad industrial adoption.
@@
With the advent of deep learning, orthographic projection reasoning has typically been cast as a standard 3D reconstruction pipeline: rasterized drawings are processed by models that output B-Rep structures [1, 2], point clouds [3, 4], or meshes [5]. However, two critical shortcomings remain. First, industrial drawings allow virtually zero tolerance for missing or erroneous components: pixel-level discretization errors in reconstruction can propagate into the 3D model, causing failures in downstream manufacturing and simulation. Second, practical CAD workflows demand parametric editability and precise semantic alignment—capabilities that common 3D representations lack. Although some methods generate CAD command sequences to restore editability, they still overlook the nuanced semantics of orthographic views and thus fail to map 2D drawing features to their intended 3D counterparts reliably.
@@
Recent works leverage the modality alignment and semantic reasoning strengths of vision–language models (VLMs) [6, 7, 8, 9, 10, 11], applying supervised fine-tuning (SFT) to address CAD challenges[12, 13, 14]. However, orthographic projection reasoning demands genuine inferential ability, and task-specific SFT often leads to memorization and out-of-distribution (OOD) degradation rather than deeper reasoning [15]. Inspired by DeepSeek R1-Zero’s Group Relative Policy Optimization (GRPO) [16], we propose Curriculum-driven Reinforcement Fine-tuning for CAD (CReFT-CAD),
@@
CReFT-CAD is a versatile framework that assists engineers in achieving orthographic projection reasoning throughout the entire lifecycle of industrial design, from design and manufacturing to simulation, in an interactive manner, as shown in Fig. 1. CReFT-CAD is a two-stage paradigm that combines reinforcement learning and supervised tuning. Curriculum-driven Reinforcement Fine-tuning: We introduce a difficulty-aware reward mechanism that incrementally escalates task complexity, gradually exposing the model to more complex reasoning challenges and promoting stable policy optimization. Supervised Post-tuning: Building on the reinforced model, we further refine its instruction-following and reasoning capabilities via multi-task SFT. Our qualitative and quantitative evaluations demonstrate that CReFT-CAD delivers robust orthographic projection reasoning and strong generalization in OOD scenarios. Another major obstacle in orthographic projection reasoning has been the lack of open-source datasets with high-fidelity annotations. To fill this gap, we introduce TriView2CAD, the first large-scale benchmark tailored for industrial CAD pipelines. Leveraging real-world design archives, we employ a constraint-driven synthesis pipeline to generate 200,000 synthetic and 3,000 real-world three-view sets, each annotated with precise, one-to-one dimension labels tied to their corresponding geometric primitives. This design enables models to extract quantitative measurements from rasterized drawings and enforce exact geometric constraints for simulation-ready 3D reconstruction.
@@ 2.1 Orthographic projection reasoning for CAD
Understanding and reasoning with orthographic projection is a fundamental and longstanding challenge in CAD-related tasks [17, 18, 19]. With the development of 3D reconstruction methods, existing orthographic projection reasoning methods have largely focused on standard 3D reconstruction pipelines. SPARE3D [20] introduces a dataset designed to evaluate the spatial reasoning capabilities of AI systems through 2D line drawings of 3D objects. Contrastive-SPARE3D [21] propose a self-supervised binary classification network that helps to learn 3D object line drawings representations that are detail-sensitive and view-invariant. PlankAssembly [22] introduces a transformer-based sequence generation model that learns flexible mappings between inputs and outputs. IsoTGAN [23] introduces a novel Gaussian-enhanced Euclidean attention mechanism and a geometric constraint loss function to enhance local image features further. GaussianCAD [24] employs a custom sparse-view 3D reconstruction method, eliminating the reliance on vector CAD sketches and 3D real-world data.
@@ 2.2 Applications of LVLMs in CAD-Related tasks
Recent work on applying vision–language models (VLMs) to CAD tasks can be grouped into three main categories: 1) Direct Adaptation of Pretrained VLMs. CAD-Recode [25] leverages a VLM to translate raw point-cloud inputs into executable Python code. Both LEAM [26] and LLM4CAD [27] generate CAD models by jointly leveraging textual descriptions and image inputs. CAD-Assistant [28] introduces a novel approach that leverages VLMs as planners, where the model interacts with Python APIs to accomplish various CAD tasks. 2) Supervised Fine-Tuning (SFT). CAD2Program [12] generates 3D parametric models from 2D CAD drawings. CAD-MLLM [29] enables parametric CAD modeling from text, images, and point clouds, and proposes the Omni-CAD dataset. 3) Reinforcement Learning–Based Tunings. RLCAD [30] proposes a reinforcement learning environment to generate complex CAD models. CADCrafter [31] presents an image-to-CAD model generation framework, fine-tuned via Direct Preference Optimization (DPO) to translate input images directly into executable CAD representations.
@@ 3 TriView2CAD: Benchmarking orthographic projection reasoning
Lack of precise dimensional annotations. Existing orthographic projection benchmarks provide only rasterized drawings without precise dimension annotations. It causes models to prioritize relative scale over exact measurements, and pixel-level errors in raster images can easily propagate into the reconstructed 3D model, resulting in inaccuracies. Additionally, the limited editability of reconstructed 3D models precludes corrective adjustments during design and simulation.
Lack of explicit logical reasoning tasks. No current dataset evaluates a model’s inferential capabilities within or across orthographic views. Consequently, models trained on these benchmarks learn pixel-level reconstruction but fail to handle real-world challenges such as inferring omitted annotations or exploiting structural symmetries.
Lack of essential CAD modalities. Most datasets include only one or two modalities—typically raster images and, at best, a single 3D format (meshes or point clouds). This limits coverage omits critical data representations (e.g., parameter tables, vector CAD files, executable commands, STEP/B-Rep) needed for end-to-end CAD design, manufacturing, and simulation pipelines.
@@ 3.1 Dataset construction
In this work, we focus on prefabricated bridge piers due to their inherently modular composition of repeated, standardized components, which yields a compact yet expressive parameter space for dataset synthesis. Moreover, prefabricated piers are ubiquitous in modern infrastructure and supported by extensive archives of CAD drawings and manufacturing documentation. TriView2CAD comprises 200,000 synthetic samples and 3,000 real-world samples. The synthetic data are split 80/20 into training and test sets, while, due to the scarcity of real-world engineering drawings, all 3,000 real-world samples are reserved exclusively for testing. The synthetic test set is used to evaluate in-domain performance. As shown in Fig. 3, real-world images typically exhibit more complex structures, primarily due to redundant annotation lines causing occlusions, overlap with other components, and the inclusion of non-numeric annotations, serving to assess out-of-distribution generalization in authentic CAD scenarios. To mirror end-to-end industrial workflows, TriView2CAD provides six interconnected data modalities per sample—structured parameter tables (JSON), vector CAD drawings (DXF), raster images (PNG), executable modeling scripts, and two standard 3D formats (STEP and B-Rep). This comprehensive suite supports tasks from early design to downstream manufacturing and simulation. As shown in Fig. 5, we employ a constraint-driven synthesis pipeline. The detailed steps are as follows:
@@
Step 1: Constraint-Guided Parameter Sampling We first analyze real-world CAD drawings to define a 15-dimensional parameter space subject to two constraint classes. Intra-View Constraints enforce both topological closure and physical validity within each projection: every component must form a gap- and overlap-free contour, and paired dimensions must satisfy domain-specific engineering constraints (e.g., “Cross-Bridge Pier Spacing” < “Cap Beam Cross-Bridge Dimension”). Inter-View Constraints ensure cross-projection consistency by requiring every measurement annotated in one view to exactly match its counterpart (height, width, and depth) in the other views. By embedding these constraints into our sampling pipeline, we guarantee that every synthetic design is geometrically coherent and adheres to real-world engineering practices.
Step 2: 2D Drawing and Raster Image Generation We employ the ezdxf library to translate each sampled parameter vector into a vectorized 2D CAD drawing (DXF), mapping numeric dimensions to primitives such as lines, circles, and arcs to form coherent orthographic projections. We leverage ezdxf’s layers-separating function to partition primitives into semantic layers, thereby enhancing the editability of the resulting CAD drawings. Each DXF file is then loaded into FreeCAD [32], where we capture high-resolution screenshots to produce three orthographic views (front, top, and side).
Step 3: 3D Model Generation We employ the FreeCAD Python API to reconstruct 3D model from its parameter vector. Mirroring our 2D pipeline, we decompose the model into individual entities and generate scripted commands that accurately position each primitive according to the intra- and inter-view constraints established in Step 1. By executing these scripts, we automatically produce industry-standard STEP and B-Rep files, closing the loop from parameter sampling to editable, simulation-ready CAD geometry.
@@ 3.2 Evaluation of TriView2CAD
We benchmark seven leading vision–language models on TriView2CAD to perform high-precision orthographic projection reasoning. The evaluation results are presented in Sec. 3.2. Reflecting real-world CAD workflows, we define three complementary evaluation tasks to rigorously test a model’s ability to extract precise dimension annotations from rasterized orthographic projection and infer the underlying geometric relationships. (i) Dimension recognition and pairing Models identify each annotated dimension in a single orthographic view and map it to its corresponding geometric feature. (ii) Primitive counting Models count instances of specified CAD elements (e.g., number of pier columns), testing their ability to parse structural composition. (iii) Composite Parameter Computation Models compute engineering-critical derived quantities—for example, the Cross-Bridge Pier Spacing is defined as the sum of the Pier Column Cross-Bridge Dimension and the Pile Spacing. Overall, we define a 15-dimensional parameter space comprising 6 recognition parameters, 3 counting parameters, and 6 composite calculation parameters. During evaluation, we treat each individual parameter as a separate prediction. The overall accuracy is then computed as the total number of correctly predicted parameters divided by the total number of parameters across all test samples. This evaluation suite exposes each model’s strengths and weaknesses in handling the tightly coupled semantic and quantitative demands of orthographic projection CAD reasoning.
@@ 4 Methodology
Inspired by the success of reinforcement learning–based tuning methods [33, 34], we present Progressive Reinforcement Fine-tuning for CAD (CReFT-CAD), a orthographic-projection reasoning framework. CReFT-CAD combines a ViT-based visual encoder with Qwen2.5 [35]. First, Progressive Reinforcement Fine-tuning employs a curriculum of three progressively harder reasoning tasks under a difficulty-aware reward scheme, enabling the model to move beyond rote pattern matching toward robust reasoning while improving OOD resilience. Second, Supervised Post-tuning refines the model’s instruction-following to support interactive, real-world CAD queries.
4.1 Curriculum-driven reinforcement fine-tuning
@@
Orthographic projection reasoning is a complex reasoning task that encompasses not only dimension recognition and pairing but also primitive counting and more intricate composite parameter computation tasks. Building on GRPO, we propose a Curriculum-driven Reinforcement Fine-tuning (CReFT) strategy, which incrementally exposes the model to progressively more complex training tasks to enhance its reasoning and generalization capabilities. Our tasks design mirrors the core steps of expert engineers in verifying CAD drawings, incorporating dichotomous choice task, multiple choice task, and parameterization tasks based on chain-of-thought (CoT). For each task, we design a tailored, efficient reward function to ensure the model can steadily and stably optimize its reasoning performance.
@@
Figure 4: Diagram design of three tasks in curriculum-driven reinforcement fine-tuning.
@@
Data engine. We leverage 160,000 image–text instruction pairs using the synthetic training set of TriView2CAD, where each prompt includes a given orthographic projection and all 15 parameter key-value pairs, as shown in Fig. 4. To ensure a balanced training signal across all three tasks, 50% of the training responses are entirely correct, while the remaining 50% contain errors of varying magnitude. Specifically, for the dichotomous choice task, we ensure that each negative response contains $n$ erroneous parameter values, where $n$ follows a uniform distribution $N\sim\text{Uniform}(a,b)$, where $a=1$ and $b=15$.
@@
For the multiple choice task, $p$ parameter values are masked in each image-text instruction pair. The masking mechanism ensures the diversity of options. This approach prevents the model from simply deriving the answer through parameterization, bypassing the complex multiple choice reasoning required in the task. Instead, it forces the model to engage in decision-making actively, considering multiple potential solutions and selecting the correct responses based on a deeper understanding. For each instruction pair, unmasked values are correct, while for the incorrect parameter value lists, all unmasked values, except for randomly selected $q$ erroneous lists. Both the masked values $p$ and the number of incorrect values $q$ follow a normal distribution, introducing a smooth and gradual variation in the difficulty of the learning task. This variation helps the model to learn in a more structured and progressive manner.
@@
For the parameterization task, we construct image-text instruction pairs based on Chain-of-Thought (CoT) reasoning. Specifically, as for Composite Parameter Computation, expert knowledge guides the model through a step-by-step reasoning guidance to arrive at the correct output. The step-by-step workflow of reasoning guidance is as follow: 1) recognize parameters required for the computation of the composite parameter; 2) provide the calculation formula for the composite parameter; 3) generate the predicted result based on the formula and the factor parameters
@@
Training strategy. Given that our task involves highly customized inputs and outputs rather than a general-purpose problem, directly applying a pretrained VLM for reasoning tasks and generating required outputs proves ineffective. As a result, the model cannot effectively execute the reinforcement learning reward mechanism, as the reward will always be 0. To address this, prior to executing each training task, we pre-sample a subset of image-text instruction pairs to "warm up" the model, facilitating the initialization of the reinforcement learning (RL) reward mechanism.

### v2

@@ Abstract
Computer-Aided Design (CAD) is pivotal in industrial manufacturing, with orthographic projection reasoning foundational to its entire workflow—encompassing design, manufacturing, and simulation. However, prevailing deep-learning approaches employ standard 3D reconstruction pipelines as an alternative, which often introduce imprecise dimensions and limit the parametric editability required for CAD workflows. Recently, some researchers adopt vision–language models (VLMs), particularly supervised fine-tuning (SFT), to tackle CAD-related challenges. SFT shows promise but often devolves into pattern memorization, resulting in poor out-of-distribution (OOD) performance on complex reasoning tasks. To tackle these limitations, we introduce CReFT-CAD, a two-stage fine-tuning paradigm: first, a curriculum-driven reinforcement learning stage with difficulty-aware rewards to steadily build reasoning abilities; second, supervised post-tuning to refine instruction following and semantic extraction. Complementing this, we release TriView2CAD, the first large-scale, open-source benchmark for orthographic projection reasoning, comprising 200,000 synthetic and 3,000 real-world orthographic projections with precise dimensional annotations and six interoperable data modalities. Benchmarking leading VLMs on orthographic projection reasoning, we show that CReFT-CAD significantly improves reasoning accuracy and OOD generalizability in real-world scenarios, providing valuable insights to advance CAD reasoning research. The code and adopted datasets are available at https://github.com/KeNiu042/CReFT-CAD.
@@ 1 Introduction
Computer‐Aided Design (CAD) is now integral to industrial product development, underpinning design, manufacturing, and simulation workflows. In the design phase, engineers use CAD drawings or rasterized orthographic projections for their precision and facile editability. During manufacturing, these drawings are converted into constraint‐based parameter tables; for simulation, they yield boundary‐representation (B‐Rep) data or textual geometry descriptions. A truly user‐centric pipeline thus requires accurate semantic parsing of orthographic projections: automated extraction of parameter tables enables direct generation of 3D models meeting stringent manufacturing and simulation standards. By contrast, existing reverse‐engineering methods rely on expensive 2D/3D scanning hardware and labor‐intensive post-processing, greatly restricting scalability and widespread industrial adoption.
@@
With the advent of deep learning, orthographic projection reasoning has typically been framed as a standard 3D reconstruction pipeline: models process rasterized drawings to output B-Rep structures [1, 2], point clouds [3, 4], or meshes [5]. However, this paradigm suffers from two critical limitations. First, industrial drawings demand near-zero tolerance for missing or erroneous components: pixel-level discretization errors in reconstruction can propagate to 3D models, leading to failures in downstream manufacturing and simulation. Second, practical CAD workflows require parametric editability and precise semantic alignment—key capabilities that common 3D representations lack. While some methods generate CAD command sequences to recover editability, they still overlook the nuanced semantics of orthographic views, thus failing to reliably map 2D drawing features to their intended 3D counterparts.
@@
Recent works leverage the modality alignment and semantic reasoning capabilities of vision–language models (VLMs) [6, 7, 8, 9, 10, 11, 12], addressing CAD challenges via supervised fine-tuning (SFT) [13, 14, 15]. However, orthographic projection reasoning requires genuine inferential ability, and task-specific SFT often induces pattern memorization and out-of-distribution (OOD) performance degradation rather than fostering deeper reasoning [16]. Inspired by DeepSeek R1-Zero’s Group Relative Policy Optimization (GRPO) [17], we propose Curriculum-driven Reinforcement Fine-tuning for CAD (CReFT-CAD).
@@
CReFT-CAD is a versatile framework enabling interactive orthographic projection reasoning across the full lifecycle of industrial design—from design and manufacturing to simulation (see Fig. 1). It consists of a two-stage paradigm integrating reinforcement learning and supervised tuning: the first stage is Curriculum-driven Reinforcement Fine-tuning, where we introduce a difficulty-aware reward mechanism that incrementally increases task complexity, gradually exposing the model to increasingly complex reasoning challenges to promote stable policy optimization; the second stage is Supervised Post-tuning, where we build on the reinforced model and further refine its instruction-following and reasoning capabilities via multi-task SFT. Our qualitative and quantitative evaluations show that CReFT-CAD achieves robust orthographic projection reasoning and strong generalization in OOD scenarios. Another key barrier in orthographic projection reasoning is the scarcity of open-source datasets with high-fidelity annotations. To address this, we present TriView2CAD, the first large-scale benchmark tailored to industrial CAD pipelines. Leveraging real-world design archives, we use a constraint-driven synthesis pipeline to generate 200,000 synthetic and 3,000 real-world three-view sets, each annotated with precise, one-to-one dimensional labels linked to their corresponding geometric primitives. This design allows models to extract quantitative measurements from rasterized drawings and enforce exact geometric constraints for 3D reconstructions ready for simulation.
@@ 2.1 Orthographic projection reasoning for CAD
Understanding and reasoning with orthographic projection is a fundamental and longstanding challenge in CAD-related tasks [18, 19, 20]. With the development of 3D reconstruction methods, existing methods for orthographic projection reasoning have largely centered on standard 3D reconstruction pipelines. SPARE3D [21] introduces a dataset for evaluating the spatial reasoning capabilities of AI systems via 2D line drawings of 3D objects. Contrastive-SPARE3D [22] proposes a self-supervised binary classification network that helps learn 3D object line drawing representations that are detail-sensitive and view-invariant. PlankAssembly [23] presents a transformer-based sequence generation model that learns flexible input-output mappings. IsoTGAN [24] introduces a novel Gaussian-enhanced Euclidean attention mechanism and a geometric constraint loss function to further enhance local image features. GaussianCAD [25] employs a custom sparse-view 3D reconstruction method, removing reliance on vector CAD sketches and real-world 3D data.
@@ 2.2 Applications of LVLMs in CAD-Related tasks
Recent efforts applying vision–language models (VLMs) to CAD tasks fall into three main categories:1) Direct Adaptation of Pretrained VLMs. CAD-Recode [26] leverages a VLM to translate raw point cloud inputs into executable Python code.Both LEAM [27] and LLM4CAD [28] generate CAD models by jointly leveraging text descriptions and image inputs. CAD-Assistant [29] introduces a novel approach using VLMs as planners, where the model interacts with Python APIs to accomplish diverse CAD tasks. 2) Supervised Fine-Tuning (SFT).CAD2Program [13] generates 3D parametric models from 2D CAD drawings. CAD-MLLM [30] enables parametric CAD modeling from text, images, and point clouds, and introduces the Omni-CAD dataset. 3) Reinforcement Learning–based Tunings.RLCAD [31] introduces a reinforcement learning environment for generating complex CAD models. CADCrafter [32] presents an image-to-CAD generation framework, fine-tuned via Direct Preference Optimization (DPO) to directly translate input images into executable CAD representations.
@@ 3 TriView2CAD: Benchmarking orthographic projection reasoning
Lack of precise dimensional annotations. Existing orthographic projection benchmarks provide only rasterized drawings without precise dimension annotations. This causes models to prioritize relative scale over exact measurements, and pixel-level errors in rasterized images can easily propagate into the reconstructed 3D model, resulting in inaccuracies. Additionally, the limited editability of reconstructed 3D models prevents corrective adjustments during design and simulation.
Lack of explicit logical reasoning tasks. Existing datasets do not evaluate a model’s inferential capabilities within or across orthographic views. Consequently, models trained on these benchmarks learn pixel-level reconstruction skills but fail to handle real-world challenges such as inferring omitted annotations or leveraging structural symmetries.
Lack of essential CAD modalities. Most datasets contain only one or two modalities—typically raster images and, at best, a single 3D format (meshes or point clouds). This limited coverage omits critical data representations (e.g., parameter tables, vector CAD files, executable commands, STEP/B-Rep) required for end-to-end CAD design, manufacturing, and simulation pipelines.
@@ 3.1 Dataset construction
In this work, we focus on prefabricated bridge piers owing to their inherently modular structure—composed of repeated, standardized components—that enables a compact yet expressive parameter space for dataset synthesis. Moreover, prefabricated piers are ubiquitous in modern infrastructure and backed by extensive CAD drawing and manufacturing documentation archives. TriView2CAD comprises 200,000 synthetic samples and 3,000 real-world samples. The synthetic data are split 80/20 into training and test sets; due to the scarcity of real-world engineering drawings, all 3,000 real-world samples are reserved exclusively for testing. The synthetic test set evaluates in-domain performance. As shown in Fig. 3, real-world images typically feature more complex structures—primarily due to redundant annotation lines (causing occlusions and overlaps with other components) and the inclusion of non-numeric annotations. These samples thus assess out-of-distribution (OOD) generalization in authentic CAD scenarios. To align with end-to-end industrial workflows, TriView2CAD provides six interconnected data modalities per sample: structured parameter tables (JSON), vector CAD drawings (DXF), raster images (PNG), executable modeling scripts, and two standard 3D formats (STEP and B-Rep). This comprehensive modality suite supports tasks spanning early design to downstream manufacturing and simulation. As shown in Fig. 5, we adopt a constraint-driven synthesis pipeline, with detailed steps outlined below:
@@
Step 1: Constraint-Guided Parameter Sampling We first analyze real-world CAD drawings to define a 15-dimensional parameter space, governed by two constraint classes. Intra-View Constraints enforce topological closure and physical validity within each projection: every component must form gap- and overlap-free contours, and paired dimensions adhere to domain-specific engineering constraints (e.g., “Cross-Bridge Pier Spacing” < “Cap Beam Cross-Bridge Dimension”). Inter-View Constraints ensure cross-projection consistency by requiring measurements annotated in one view to exactly match their counterparts (height, width, depth) in other views. By embedding these constraints into the sampling pipeline, we ensure all synthetic designs are geometrically coherent and compliant with real-world engineering standards.
Step 2: 2D Drawing and Raster Image Generation We use the ezdxf library to convert each sampled parameter vector into a vectorized 2D CAD drawing (DXF format), mapping numeric dimensions to geometric primitives (e.g., lines, circles, arcs) to form coherent orthographic projections. We utilize ezdxf’s layer separation functionality to partition primitives into semantic layers, enhancing the editability of the resulting CAD drawings. Each DXF file is then imported into FreeCAD [33], where high-resolution screenshots are captured to generate three orthographic views (front, top, side).
Step 3: 3D Model Generation We utilize the FreeCAD Python API to reconstruct a 3D model from its corresponding parameter vector. Consistent with our 2D generation pipeline, we decompose the model into discrete entities and generate scripted commands to precisely position each primitive in compliance with the intra- and inter-view constraints defined in Step 1. Executing these scripts automatically produces industry-standard STEP and B-Rep files, closing the loop from parameter sampling to editable, simulation-ready CAD geometries.
@@ 3.2 Evaluation of TriView2CAD
We benchmark seven leading vision–language models on TriView2CAD to assess their high-precision orthographic projection reasoning capabilities. The evaluation results are presented in Sec. 3.2. Aligning with real-world CAD workflows, we design three complementary evaluation tasks to rigorously test a model’s ability to extract precise dimension annotations from rasterized orthographic projections and infer the underlying geometric relationships: (i) Dimension recognition and pairing Models identify each annotated dimension in a single orthographic view and map it to its associated geometric feature. (ii) Primitive counting Models count instances of specified CAD elements (e.g., number of pier columns), assessing their ability to parse structural composition. (iii) Composite Parameter Computation Models compute engineering-critical derived quantities—for example, Cross-Bridge Pier Spacing is calculated as the sum of Pier Column Cross-Bridge Dimension and Pile Spacing. Collectively, we define a 15-dimensional parameter space consisting of 6 recognition parameters, 3 counting parameters, and 6 composite calculation parameters. In evaluation, each parameter is treated as an independent prediction target. Overall accuracy is computed as the total number of correctly predicted parameters divided by the total number of parameters across all test samples. This evaluation suite reveals the strengths and weaknesses of each model in addressing the tightly coupled semantic and quantitative requirements of orthographic projection CAD reasoning.
@@ 4 Methodology
Inspired by the success of reinforcement learning–based tuning methods [34, 35], we present Curriculum-driven Reinforcement Fine-tuning for CAD (CReFT-CAD), an orthographic-projection reasoning framework. CReFT-CAD integrates a ViT-based visual encoder with Qwen2.5 [36]. First, the framework adopts a curriculum of three progressively challenging reasoning tasks paired with a difficulty-aware reward mechanism, empowering the model to transcend rote pattern matching and develop robust reasoning capabilities while enhancing out-of-distribution (OOD) generalization. Second, Supervised Post-tuning refines the model’s instruction-following abilities to accommodate interactive, real-world CAD queries.
4.1 Curriculum‐driven reinforcement fine-tuning
@@
Orthographic projection reasoning is a sophisticated reasoning task that involves not only dimension recognition and pairing but also primitive counting and more complex composite parameter computation. Building upon GRPO, we propose a Curriculum-driven Reinforcement Fine-tuning (CReFT) strategy, which incrementally exposes the model to increasingly complex training tasks to enhance its reasoning and generalization capabilities. Our task design aligns with the core steps expert engineers follow when verifying CAD drawings, incorporating dichotomous choice tasks, multiple-choice tasks, and chain-of-thought (CoT)-based parameterization tasks. For each task, we design a tailored, efficient reward function to ensure the model can steadily optimize its reasoning capabilities.
@@
Figure 4: Diagram design of three tasks in curriculum‐driven reinforcement fine-tuning.
@@
Data engine. We utilize 160,000 image–text instruction pairs from the synthetic training set of TriView2CAD, where each prompt includes an orthographic projection and all 15 parameter key-value pairs (see Fig. 4).To ensure a balanced training signal across the three tasks, 50% of the training responses are fully correct, while the remaining 50% contain errors of varying degrees. Specifically, for dichotomous choice tasks, each negative response includes $n$ erroneous parameter values, where $n$ follows a uniform distribution $N\sim\text{Uniform}(a,b)$, where $a=1$ and $b=15$.
@@
For multiple-choice tasks, $p$ parameter values are masked out in each image-text instruction pair. The masking mechanism is designed to ensure the diversity of distractor options, preventing the model from merely deriving answers through parameter matching—thereby avoiding circumvention of the complex reasoning required for the task. Instead, it compels the model to actively engage in decision-making, evaluating multiple potential solutions and selecting the correct one based on a deeper geometric understanding. For each instruction pair, unmasked values are correct; for incorrect parameter value lists, all unmasked values are correct except for those in $q$ randomly selected erroneous entries. Both the number of masked values $p$ and the number of erroneous entries $q$ follow normal distributions, introducing smooth, gradual variations in task difficulty that enable the model to learn in a more structured, progressive fashion.
@@
For parameterization tasks, we construct image-text instruction pairs grounded in Chain-of-Thought (CoT) reasoning. Specifically, for Composite Parameter Computation tasks, expert knowledge is embedded into step-by-step reasoning prompts to guide the model toward correct outputs. The step-by-step reasoning workflow is as follows: 1) identify the parameters required to compute the composite parameter; 2) specify the calculation formula for the composite parameter; 3) compute the predicted result using the formula and relevant factor parameters.
@@
Training strategy. Given that our task involves highly specialized inputs and outputs (as opposed to general-purpose problems), directly applying a pretrained VLM to perform reasoning and generate the required outputs is ineffective. Consequently, the reinforcement learning (RL) reward mechanism cannot function properly—with rewards consistently remaining zero. To mitigate this, we pre-sample a subset of image-text instruction pairs to "warm up" the model before initiating each training task, enabling proper initialization of the RL reward mechanism.